OpenStudy (anonymous):

1/7x^2+1/7x-2/7. what are all real zeros

OpenStudy (anonymous):

please clarify, is this what you want?\[\frac{ 1 }{ 7 }*x^{2} + \frac{ 1 }{ 7 } *x - \frac{2}{7}\]

OpenStudy (anonymous):

yes it is. thank you im so confused because of the fractions

OpenStudy (anonymous):

okay, as this is a parabola you can use the formula and do \[\frac{ -b + \sqrt{b^2 - 4 a c} }{ 2a }\] your a will be 1/7 the b 1/7 and c -2/7

OpenStudy (anonymous):

You can see the parabola and more infos about it in here: http://www.wolframalpha.com/input/?i=%281%2F7%29+x%5E2+%2B+%281%2F7%29+x+-+2%2F7 If you have any more doubts just ask

OpenStudy (anonymous):

thank you so much man